### v2

@@ 7.3.2 Finite-memory Nash equilibria in co-Büchi games
For each $i\in\llbracket{}n\rrbracket{}$, we define a Mealy machine $\mathcal{M}_{i}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ as follows. We define $M$ as the set $\{v_{\ell}\in V\mid 0\leq\ell\leq r^{\star}\}\cup\{(\mathcal{P}_{i},1)\mid i\in I\}\cup\{\mathcal{P}_{i}\mid i\in I\}$ and $m_{\mathsf{init}}=v_{0}$. The memory states that are vertices correspond to the first phase in the sketch of the main text (enforcing the history $\pi{{r^{\star}}}$), and the others to the second phase (following a strategy based on a trivial decomposition). The memory bounds claimed for finite arenas hold by simplicity of $\pi{{r^{\star}}}$.
The update function $\mathsf{up}$ is defined as follows. For all $\ell<{r^{\star}}$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$. We postpone the definition of $\mathsf{up}(v_{r^{\star}},v_{r^{\star}})$. For all $\ell\leq{r^{\star}}$ and all $v\neq v_{\ell}$, if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$ (i.e., $v\in V_{I}$), we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$, and otherwise the update is arbitrary. For all $i\in I$ and $v\in V$ occurring in $\pi{{r^{\star}}}$, we let $\mathsf{up}((\mathcal{P}_{i},1),v)=(\mathcal{P}_{i^{\prime}},1)$ where $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$. For all $i\in I$ and all $v\in V$ that do not occur in $\pi{{r^{\star}}}$, we let $\mathsf{up}((\mathcal{P}_{i},1),v)=\mathcal{P}_{i}$. We let $\mathsf{up}(v_{r^{\star}},v_{r^{\star}})=\mathsf{up}((\mathcal{P}_{i},1),v_{r^{\star}})$ for some $i\in I$ (the definition does not depend on $i$ because $v_{r^{\star}}$ occurs in $\pi{{r^{\star}}}$). Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$. We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq{r^{\star}}$. If $v=v_{\ell}$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$ if $\ell\neq{r^{\star}}$ and let $\mathsf{nxt}_{i}(v_{r^{\star}},v_{r^{\star}})$ be the second vertex of $\pi{{r^{\star}}}$ (i.e., the only vertex that follows $v_{{r^{\star}}}$ in $\pi{r^{\star}}$) if $\ell={r^{\star}}$. Now, assume that $v\neq v_{\ell}$. If $\ell\geq 1$ and $v_{\ell-1}\in V_{I}$, we let $i^{\prime}\in I$ such that $v\in V_{i^{\prime}}$ and set $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$. We let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases. We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},1)$. Fix $i^{\prime}\in I$. If $v$ occurs in $\pi{r^{\star}}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},1),v)$ be the vertex following $v$ in $\pi{r^{\star}}$ (this vertex is well-defined as $\pi$ is a simple play or a simple lasso). Otherwise, if $v$ does not occur in $\pi{r^{\star}}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},1),v)=\tau_{-i^{\prime}}(v)$. Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and otherwise we let it be arbitrary.
@@
Second, we assume that we can write $\pi^{\prime}$ as a combination $\pi{r^{\star}}\cdot\pi^{\prime\prime}$ and that all states that occur in $\pi^{\prime\prime}$ also occur in $\pi{r^{\star}}$. By definition of the co-Büchi objective, $\mathsf{SatPl}_{\mathcal{G}}(\pi)\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$. Furthermore, as $\pi$ was obtained through Lemma 7.13ii implies that $\mathsf{SatPl}_{\mathcal{G}}(\pi)\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$, and, in particular, $\pi^{\prime}\in\mathsf{B\ddot{u}chi}(T_{i})$.
@@
We establish the existence of memoryless uniformly optimal strategies of $\mathcal{P}_{1}$ in $\mathcal{G}$ with the following steps. First, we argue the existence of optimal strategies for $\mathcal{P}_{1}$. Second, we construct a shortest-path game by removing edges from $\mathcal{A}$ without introducing deadlocks. We then show that values coincide in this new game and the original game. Finally, we establish that memoryless uniformly winning reachability strategies of the new game are optimal in both this new game and the original shortest-path game. See 3.5
@@
First, we establish that there exists optimal strategies in $\mathcal{G}$ from each vertex. Let $v\in V$. By Lemma 3.3, $\mathsf{val}(v)$ is defined and its proof implies that we have $\mathsf{val}(v)\in\overline{\mathbb{N}}$. First, assume that $\mathsf{val}(v)$ is finite, i.e., $\mathsf{val}(v)\in\mathbb{N}$. By definition of the value, $\mathcal{P}_{1}$ has a strategy that ensures $\mathsf{val}(v)+\frac{1}{2}$ from $v$. This strategy is optimal due to the codomain of the shortest-path cost. Second, assume that $\mathsf{val}(v)=+\infty$. In this case, all strategies of $\mathcal{P}_{1}$ are optimal.
Second, we prove that for all $v\in V_{1}\setminus T$, there exists $v^{\prime}\in\mathsf{Succ}_{E}(v)$ such that $\mathsf{val}(v)=\mathsf{val}(v^{\prime})+w(v,v^{\prime})$. Let $v\in V_{1}\setminus T$. If $\mathcal{P}_{1}$ moves from $v$ to $v^{\prime}\in\mathsf{Succ}_{E}(v)$, $\mathcal{P}_{1}$ can ensure $\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ at best. It follows that $\mathsf{val}(v)=\min\{\mathsf{val}(v^{\prime})+w(v,v^{\prime})\mid v^{\prime}\in\mathsf{Succ}_{E}(v)\}$ (this minimum is well-defined because $\overline{\mathbb{N}}$ is well-ordered). This implies the claim.
Third, we claim that for all $v\in V_{2}\setminus T$, we have $\mathsf{val}(v)\geq\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ for all $v^{\prime}\in\mathsf{Succ}_{E}(v)$. Let $v\in V_{2}\setminus T$ and $v^{\prime}\in\mathsf{Succ}_{E}(v)$. Assume first that $\mathsf{val}(v^{\prime})$ is finite, $\mathcal{P}_{2}$ has a strategy ensuring $\mathsf{val}(v)-\frac{1}{2}$ from $v$, which is optimal. It follows that $\mathcal{P}_{2}$ can ensure $\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ from $v$ by moving to $v^{\prime}$ from $v$ and playing optimally from there, which implies the desired inequality. Assume now that $\mathsf{val}(v^{\prime})$ is infinite. Then for all $\alpha\in\mathbb{N}$, $\mathcal{P}_{2}$ has a strategy ensuring $\alpha$ from $v^{\prime}$. We conclude, similarly to the previous case, that $\mathsf{val}(v)$ is infinite and therefore satisfies the inequality.
Fourth, we remove edges of $\mathcal{A}$ to derive a game in which values are unchanged with respect to $\mathcal{G}$. Let $E^{\prime}$ be obtained by removing from $E$ all edges $(v,v^{\prime})$ with $v\in V_{1}\setminus T$ and $\mathsf{val}(v)\neq\mathsf{val}(v^{\prime})+w(v,v^{\prime})$. We let $\mathcal{A}^{\prime}=((V_{1},V_{2}),E^{\prime})$ and $\mathcal{G}^{\prime}=(\mathcal{A}^{\prime},\mathsf{SPath}^{T}_{w})$. We remark that there are no deadlocks in $\mathcal{A}^{\prime}$ by our second point. For any $v\in V$, let $\mathsf{val}^{\prime}(v)$ denote its value in $\mathcal{G}^{\prime}$. We claim that
@@
For i, we observe that for all $v\in V$, $\mathsf{val}^{\prime}(v)\geq\mathsf{val}(v)$ by definition of $\mathcal{G}^{\prime}$ ($\mathcal{P}_{1}$ has less choices than in $\mathcal{G}$). In particular, if $\mathsf{val}(v)=+\infty$, then $\mathsf{val}^{\prime}(v)=+\infty$. We show the other inequality of i by induction on $\mathsf{val}(v)$ for vertices of finite value. For the base case, let $v\in V$ such that $\mathsf{val}(v)=0$. In this case, $\mathcal{P}_{1}$ has a strategy that reaches $T$ from $v$ using edges with zero weight that moves only to vertices with zero value until a target is reached (otherwise $\mathcal{P}_{1}$ could not ensure a cost of 00 from $v$). These edges are in $E^{\prime}$. This ends the argument for the base case.
@@
Finally, we prove the claim of the theorem. Let $\sigma_{1}$ be a memoryless uniformly winning reachability strategy in the reachability game $(\mathcal{A}^{\prime},\mathsf{Reach}(T))$. It follows from ii of our fourth point that is it also a uniformly winning strategy in $(\mathcal{A},\mathsf{Reach}(T))$. It remains to show that $\sigma_{1}$ is optimal from all vertices with finite value. Let $v_{0}\in V$ such that $\mathsf{val}(v_{0})$ is finite. Let $\pi=v_{0}v_{1}\ldots$ be consistent with $\sigma_{1}$. We claim that $\mathsf{SPath}^{T}_{w}(\pi)\leq\mathsf{val}(v_{0})$. Let $r=\min\{\ell\in\mathbb{N}\mid v_{\ell}\in T\}$ (it exists by ii and the fact $\mathsf{val}(v_{0})$ is finite). We have, by choice of $E^{\prime}$ and the third claim above, that

### v3

@@ 7.3.2 Finite-memory Nash equilibria in co-Büchi games
For each $i\in\llbracket{}n\rrbracket{}$, we define a Mealy machine $\mathcal{M}_{i}=(M,m_{\mathsf{init}},\mathsf{up},\mathsf{nxt}_{i})$ as follows. We define $M$ as the set $\{v_{\ell}\in V\mid 0\leq\ell\leq r^{\star}\}\cup\{(\mathcal{P}_{i},1)\mid i\in I\}\cup\{\mathcal{P}_{i}\mid i\in I\}$ and $m_{\mathsf{init}}=v_{0}$. With respect to the informal explanation preceding the theorem (and as was the case for Büchi games), the memory states that are vertices correspond to the first phase (enforcing the history $\pi{{r^{\star}}}$), and the others to the second phase (following a strategy based on a simple decomposition). The memory bounds claimed for finite arenas hold by simplicity of $\pi{{r^{\star}}}$.
The update function $\mathsf{up}$ is defined as follows.
  1. (a)
    For all $\ell<{r^{\star}}$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$.
  2. (b)
    For all $\ell\leq{r^{\star}}$ and all $v\neq v_{\ell}$,
    • if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$, we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$;
    • otherwise the update is arbitrary.
  3. (c)
    For all $i\in I$ and $v\in V$ occurring in $\pi{{r^{\star}}}$, we let $\mathsf{up}((\mathcal{P}_{i},1),v)=(\mathcal{P}_{i^{\prime}},1)$ where $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$.
  4. (d)
    For all $i\in I$ and all $v\in V$ that do not occur in $\pi{{r^{\star}}}$, we let $\mathsf{up}((\mathcal{P}_{i},1),v)=\mathcal{P}_{i}$.
  5. (e)
    We let $\mathsf{up}(v_{r^{\star}},v_{r^{\star}})=\mathsf{up}((\mathcal{P}_{i},1),v_{r^{\star}})$ for some $i\in I$ (the choice of $i$ here does not impact the following arguments because $v_{r^{\star}}$ occurs in $\pi{{r^{\star}}}$).
  6. (f)
    Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$.
  1. (a)
    We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq{r^{\star}}$.
    • If $v=v_{\ell}$ and $\ell\neq{r^{\star}}$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$.
    • If $v=v_{\ell}$ and $\ell={r^{\star}}$, we let $\mathsf{nxt}_{i}(v_{r^{\star}},v_{r^{\star}})$ be the second vertex of $\pi{{r^{\star}}}$ (i.e., the only vertex that follows $v_{{r^{\star}}}$ in $\pi{r^{\star}}$) if $\ell={r^{\star}}$.
    • We consider two cases whenever $v\neq v_{\ell}$.
      • If $\ell\geq 1$ and there exists $i^{\prime}\in I$ such that $v\in V_{i^{\prime}}$ and $i^{\prime}\neq i$, we let and set $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$.
      • We let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases.
  2. (b)
    We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},1)$. Fix $i^{\prime}\in I$.
    • If $v$ occurs in $\pi{r^{\star}}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},1),v)$ be the vertex following $v$ in $\pi{r^{\star}}$ (this vertex is well-defined as $\pi$ is a simple play or a simple lasso).
    • If $v$ does not occur in $\pi{r^{\star}}$ and $i^{\prime}\neq i$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},1),v)=\tau_{-i^{\prime}}(v)$.
    • If $v$ does not occur in $\pi{r^{\star}}$ and $i^{\prime}=i$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},1),v)$ be arbitrary.
  3. (c)
    Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and otherwise we let it be arbitrary.
@@
Second, we assume that we can write $\pi^{\prime}$ as a combination $\pi{r^{\star}}\cdot\pi^{\prime\prime}$ and that all states that occur in $\pi^{\prime\prime}$ also occur in $\pi{r^{\star}}$. By definition of the co-Büchi objective, $\mathsf{SatPl}_{\mathcal{G}}(\pi)\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$. Furthermore, as $\pi$ was obtained through Lemma 7.14, ii implies that $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$, and, in particular, $\pi^{\prime}\in\mathsf{B\ddot{u}chi}(T_{i})$.
@@
8 Conclusion
In this paper, we have provided constructions for finite-memory NEs that are based on a relaxation of the classical punishment mechanism: players punish some deviations while others are left unpunished. We obtain finite-memory strategies as follows: instead of encoding an entire play in the memory of the players (as is done in classical approaches), we encode information regarding a decomposition of a play along with information on the latest players to have moved, to discern who to punish when necessary. We have used variants of this idea to show that in reachability, shortest-path and safety games, a memory quadratic in the number of players is sufficient to implement an NE with an outcome whose cost profile is componentwise smaller than or equal to that of some given NE, and to prove that finite-memory is always sufficient for the same problem in Büchi and co-Büchi games. The limits of the approach can be seen in Büchi and co-Büchi games with Proposition 7.3: it is no longer possible to obtain arena-independent memory upper bounds.
We discuss two natural directions for future work. A first direction would be to study the usefulness of decomposition-based approaches to refine memory bounds for (constrained) Nash equilibria in games with other types of objectives. A natural candidate objective would the parity objective, which generalises both Büchi and co-Büchi objectives. Any construction showing that finite memory suffices in parity games would yield a unified construction for Büchi games and co-Büchi games. We remark that the two-phase approach (enforce a prefix then use a decomposition-based strategy) appears to not directly extend to parity games. Consider the arena of Figure 9 and the (parity) game in which the objective of $\mathcal{P}_{1}$ is $\mathsf{B\ddot{u}chi}(\{v_{1}\})$ and the objective of $\mathcal{P}_{2}$ is $\mathsf{coB\ddot{u}chi}(\{v_{1},v_{3}\})$. The play $(v_{0}v_{1}v_{2})^{\omega}$ is arguably the simplest NE outcome from $v_{0}$ that satisfies the objective of $\mathcal{P}_{1}$. A natural candidate decomposition for the second phase of (a variant of) the two-phase mechanism would be some trivial decomposition such as $((v_{0}v_{1}v_{2})^{\omega})$ or a periodic decomposition such as $(v_{0}v_{1}v_{2}v_{0},v_{0}v_{1}v_{2}v_{0},\ldots)$. In either case, once the second phase starts and $\mathcal{P}_{1}$ uses a decomposition-based strategy, $\mathcal{P}_{2}$ has a profitable deviation by using the edge from $v_{0}$ to $v_{2}$ to avoid $v_{1}$. This suggests that a different approach may be necessary to study the more general parity games.
$v_{0}$$v_{1}$$v_{2}$$v_{3}$
Figure 9: A two-player arena (circles and squares are respectively vertices of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$) illustrating that the two-phase approach used in Büchi and co-Büchi games does not directly generalise to games with both Büchi and co-Büchi objectives: if the objective of $\mathcal{P}_{1}$ is $\mathsf{B\ddot{u}chi}(\{v_{2}\})$ and the objective of $\mathcal{P}_{2}$ is $\mathsf{coB\ddot{u}chi}(\{v_{2},v_{3}\})$, then there is no NE in which $\mathcal{P}_{1}$ wins that eventually conforms to a decomposition-based strategy with respect to $(v_{0}v_{1}v_{2})^{\omega}$.
A second direction would be to investigate the use of decomposition-based strategies to obtain results on memory requirements for equilibria besides Nash equilibria, such as secure equilibria and subgame perfect equilibria.
Acknowledgements
The author expresses their thanks to Thomas Brihaye, Aline Goeminne and Mickael Randour for fruitful discussions and their comments on a preliminary version of this paper.
@@ Acknowledgements
We establish the existence of memoryless uniformly optimal strategies of $\mathcal{P}_{1}$ in $\mathcal{G}$ by exploiting the existence of memoryless uniformly winning strategies in zero-sum reachability games (Theorem 3.1). However, we cannot directly apply this result to the reachability game $(\mathcal{A},\mathsf{Reach}(T))$ as memoryless uniformly winning strategies in this game need not be optimal in $\mathcal{G}$. Instead, we modify $\mathcal{A}$ by removing edges of $\mathcal{P}_{1}$ that are not optimal in one step to obtain an arena $\mathcal{A}^{\prime}$. We identify these edges by studying the value of vertices. We then conclude by showing that memoryless uniformly winning strategies in $(\mathcal{A}^{\prime},\mathsf{Reach}(T))$ are uniformly optimal strategies in $\mathcal{G}$.
See 3.5
@@
We recall that there exists optimal strategies of $\mathcal{P}_{1}$ in $\mathcal{G}$ from each vertex by Lemma 3.3.
First, we prove that for all $v\in V_{1}\setminus T$, there exists $v^{\prime}\in\mathsf{Succ}_{E}(v)$ such that $\mathsf{val}(v)=\mathsf{val}(v^{\prime})+w(v,v^{\prime})$. Let $v\in V_{1}\setminus T$. If $\mathcal{P}_{1}$ moves from $v$ to $v^{\prime}\in\mathsf{Succ}_{E}(v)$, $\mathcal{P}_{1}$ can ensure $\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ at best. It follows that $\mathsf{val}(v)=\min\{\mathsf{val}(v^{\prime})+w(v,v^{\prime})\mid v^{\prime}\in\mathsf{Succ}_{E}(v)\}$ (this minimum is well-defined because values are in the well-ordered set $\overline{\mathbb{N}}$). This implies the claim.
Second, we claim that for all $v\in V_{2}\setminus T$, we have $\mathsf{val}(v)\geq\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ for all $v^{\prime}\in\mathsf{Succ}_{E}(v)$. Let $v\in V_{2}\setminus T$ and $v^{\prime}\in\mathsf{Succ}_{E}(v)$. Assume first that $\mathsf{val}(v^{\prime})$ is finite, in which case $\mathcal{P}_{2}$ has an optimal strategy from $v^{\prime}$ by Lemma 3.3. It follows that $\mathcal{P}_{2}$ can ensure $\mathsf{val}(v^{\prime})+w(v,v^{\prime})$ from $v$ by moving to $v^{\prime}$ from $v$ and playing optimally from there, which implies the desired inequality. Assume now that $\mathsf{val}(v^{\prime})$ is infinite. Then for all $\alpha\in\mathbb{N}$, $\mathcal{P}_{2}$ has a strategy ensuring $\alpha$ from $v^{\prime}$. We conclude, similarly to the previous case, that $\mathsf{val}(v)$ is infinite and therefore satisfies the inequality.
Third, we remove edges of $\mathcal{A}$ to derive a game in which values are unchanged with respect to $\mathcal{G}$. Let $E^{\prime}$ be obtained by removing from $E$ all edges $(v,v^{\prime})$ with $v\in V_{1}\setminus T$ and $\mathsf{val}(v)\neq\mathsf{val}(v^{\prime})+w(v,v^{\prime})$. We let $\mathcal{A}^{\prime}=((V_{1},V_{2}),E^{\prime})$ and $\mathcal{G}^{\prime}=(\mathcal{A}^{\prime},\mathsf{SPath}^{T}_{w})$. We remark that there are no deadlocks in $\mathcal{A}^{\prime}$ by the first point above. For any $v\in V$, let $\mathsf{val}^{\prime}(v)$ denote its value in $\mathcal{G}^{\prime}$. We claim that
@@
For i, we observe that for all $v\in V$, $\mathsf{val}^{\prime}(v)\geq\mathsf{val}(v)$ by definition of $\mathcal{G}^{\prime}$ ($\mathcal{P}_{1}$ has less choices than in $\mathcal{G}$). In particular, if $\mathsf{val}(v)=+\infty$, then $\mathsf{val}^{\prime}(v)=+\infty$. We show the other inequality of i by induction on $\mathsf{val}(v)$ for vertices of finite value. For the base case, let $v\in V$ such that $\mathsf{val}(v)=0$. In this case, $\mathcal{P}_{1}$ has a strategy that reaches $T$ from $v$ using edges with zero weight that moves only to vertices with zero value until a target is reached (otherwise $\mathcal{P}_{1}$ could not ensure a cost of 0 from $v$). These edges are in $E^{\prime}$. This ends the argument for the base case.
@@
Finally, we prove the claim of the theorem. Let $\sigma_{1}$ be a memoryless uniformly winning reachability strategy in the reachability game $(\mathcal{A}^{\prime},\mathsf{Reach}(T))$. It follows from ii above that is it also a uniformly winning strategy in $(\mathcal{A},\mathsf{Reach}(T))$. It remains to show that $\sigma_{1}$ is optimal from all vertices with finite value. Let $v_{0}\in V$ such that $\mathsf{val}(v_{0})$ is finite. Let $\pi=v_{0}v_{1}\ldots$ be consistent with $\sigma_{1}$. We claim that $\mathsf{SPath}^{T}_{w}(\pi)\leq\mathsf{val}(v_{0})$. Let $r=\min\{\ell\in\mathbb{N}\mid v_{\ell}\in T\}$ (it exists by ii and the fact $\mathsf{val}(v_{0})$ is finite). We have, by choice of $E^{\prime}$ and the second claim above, that